# Electron-Phonon interaction and lattice thermal conductivity from metals to 2D Dirac crystals: a review

### v1

@@ 4.6 Summary and open challenges
It should be noted that in calculating the e-ph interaction, the Thomas–Fermi approximation, predicated on a large, nearly spherical Fermi surface and ultrashort screening lengths, breaks down in 2D Dirac crystals, with strong e-ph coupling and where the Fermi “surface” collapses to a point and the carrier density sits far below the range in which the Thomas-Fermi screening is valid. Consequently, the electronic screening length can be comparable to, or even exceed, the relevant phonon wavelength, and the static Thomas-Fermi response fails to reproduce the sharp enhancement that occurs when the phonon wave-vector approaches the nesting condition $q\simeq 2k_{F}$. It has been demonstrated that evaluating the full Lindhard polarisation for acoustic phonons yields a dielectric function with a pronounced cusp at $q=2k_{F}$ in the static limit $(\omega=0)$ and a divergence in the dynamic regime—behaviour entirely missed by Thomas-Fermi theory.[27] This analysis replaces the oversimplified Thomas-Fermi picture with a dynamic, $q$- and $\omega$-dependent Lindhard dielectric function that faithfully captures the strong, wave-vector-selective e–ph coupling characteristic of 2D Dirac crystals.
@@ 5.1 Branch dependence of e-ph lifetimes and their impact on the thermal conductivity
with bending rigidity $\kappa_{b}$ and areal mass density $\rho$. In two dimensions these dispersions imply that the phonon density of states (DOS) scales as $D_{\mathrm{LA/TA}}(\omega)\!\propto\!\omega$ for linear modes and is approximately constant for the quadratic ZA branch (contrary to a $\omega^{-1/2}$ scaling); this difference is central to the low–frequency phase space in 2D.[13, 97, 98]
@@
$\Gamma_{\nu\mathbf{q}}\equiv\bigl{(}\tau^{\mathrm{e\text{-}ph}}_{\nu\mathbf{q}}\bigr{)}^{-1}=\frac{2\pi}{\hbar}\!\sum_{mn\mathbf{k}}\bigl{|}g^{\nu}_{mn}(\mathbf{k},\mathbf{q})\bigr{|}^{2}\bigl{[}f_{n\mathbf{k}}-f_{m,\mathbf{k+q}}\bigr{]}\,\delta\bigl{(}\epsilon_{n\mathbf{k}}-\epsilon_{m,\mathbf{k+q}}+\hbar\omega_{\nu\mathbf{q}}\bigr{)},$ (35)
where the Pauli factor selects states near $E_{\mathrm{F}}$ and $g^{\nu}_{mn}$ are the DFPT/Wannier matrix elements.[48, 99] Before specializing the small $\omega$ scalings of the e-ph linewidths, it is useful to recall how the BTE integrand-the product of the phonon density of states $D_{\nu}(\omega)$, the modal heat capacity $C_{\nu}(\omega)$, the squared group velocity $v_{\nu}(\omega)]^{2}$, and the lifetime $\tau_{\nu}(\omega)$—weights each branch in suspended graphene where the first-order ZA e-ph vertex is symmetry-forbidden. Full, iterative solutions of the phonon BTE (including normal processes) predict a ZA-dominated lattice conductivity and a large gap relative to single-mode RTA. This baseline (ph–ph limited, $\sigma_{h}$ intact) is summarized in Fig. 4(a).
First, in-plane LA/TA modes. At long wavelength, graphene’s in–plane coupling is governed by an unscreened gauge–field (shear) interaction, while the scalar deformation potential is strongly screened. First–principles BTE shows that this gauge term controls electronic resistivity and the low–$q$ e–ph matrix elements for LA/TA modes.[99, 100] Near charge neutrality the available electronic phase space is small, but for doped samples $N(E_{\mathrm{F}})\!\propto\!|E_{\mathrm{F}}|$ grows linearly with $|E_{\mathrm{F}}|$ (or $k_{F}$) and enhances $\Gamma_{\mathrm{LA/TA}}$ roughly in proportion to $k_{F}$ at fixed $\omega$. Consequently, the LA/TA contribution to the e–ph–limited lattice conductivity decreases with carrier density through the $k_{F}$–dependence of $\tau_{\mathrm{LA/TA}}$, while retaining the weak qq–dependence characteristic of linear modes.[91, 92]
At long wavelength the gauge-field (shear) coupling dominates and is only weakly $q$-dependent, whereas the scalar deformation potential is strongly screened.[99, 100] In doped graphene this gives the small-$\omega$ scaling
@@
consistent with the linear-$T$ acoustic resistivity of Dirac carriers.[91, 92]
Second, flexural ZA modes and mirror symmetry. For freestanding, mirror–symmetric graphene the first–order (one–phonon) e–ph coupling to ZA modes is symmetry–forbidden; ZA scattering of electrons then proceeds via higher–order/two–phonon processes and is comparatively weak.[93] This same mirror selection rule also restricts anharmonic ph–ph scattering of ZA modes, which helps explain why ZA phonons dominate $\kappa_{\mathrm{ph}}$ in pristine suspended graphene.[13, 101] When mirror symmetry is broken by a substrate, electrostatic gating, static ripples or edges, a linear one–phonon ZA coupling appears with small-$q$ amplitude $\propto q$, giving
@@
This greatly increases the ZA damping rate and reduces the ZA share of $\kappa_{\mathrm{ph}}$. First–principle studies show this mechanism can become mobility/thermal-transport limiting in supported devices and under gating.[98, 102]
@@
with bending rigidity $\kappa_{b}$ and areal mass density $\rho$. In two dimensions these dispersions imply that the phonon density of states (DOS) scales as $D_{\mathrm{LA/TA}}(\omega)\!\propto\!\omega$ for linear modes and is approximately constant for the quadratic ZA branch (contrary to a $\omega^{-1/2}$ scaling); this difference is central to the low–frequency phase space in 2D.[97, 98, 13]
@@
showing a stronger ultraviolet weighting than LA/TA. In pristine suspended graphene the linear ZA coupling is forbidden, $\Gamma_{\mathrm{ZA}}\!\to\!0$ at one-phonon level, and ZA modes dominate $\kappa_{\mathrm{ph}}$; on substrates or under gating, the activated ZA coupling enhances ZA damping and shifts heat carriage toward in-plane modes.[98, 13, 102] It should be noted that Eqs. (41)–(42) capture robust trends. Quantitative predictions require the full mode and $\mathbf{q}$ resolved $g^{\nu}_{mn}$, realistic low-$\omega$ cutoffs (finite size/tension for ZA), and the actual Debye (or numerical) upper limit.[48, 13, 99]
@@ 5.2 Effect of tensile strain on flexural modes and the thermal conductivity
Applying a uniform uniaxial tensile strain $\epsilon$ couples the bending rigidity $\kappa_{b}$ to an in-plane membrane tension $\sigma=Y\,\epsilon$ (with $Y$ the 2D Young’s modulus). The flexural (ZA) dispersion is renormalized from purely quadratic to[98, 13]
@@
Second, higher group velocity and modified scattering. Linearization raises the ZA group velocity from $v_{\mathrm{ZA}}=\partial\omega/\partial q\!\propto\!q$ (bending) to the strain-controlled constant $v_{\epsilon}$ at small $q$. Combined with the DOS change, this alters the three-phonon phase space and lengthens ZA lifetimes in suspended graphene; within the BTE, larger $v^{2}\tau$ for ZA modes increases their contribution to $\kappa_{\mathrm{ph}}$.[13]
First-principles BTE calculations show that even modest tensile strain $(\approx 1\%)$ linearizes ZA at long wavelengths and increases the intrinsic lattice thermal conductivity of suspended graphene; in the thermodynamic limit, $\kappa_{\mathrm{ph}}$ becomes length-divergent under tension due to the altered ZA scattering landscape.[13] For finite samples, the net gain depends on size and edge scattering: when the characteristic length exceeds the dominant ZA mean-free paths (tens of microns and beyond), strain yields sizable $\kappa_{\mathrm{ph}}$ enhancements; for smaller membranes or ribbons, boundary scattering can mask the intrinsic increase, leading to weaker strain dependence.[98, 103] Overall, strain-induced flexural hardening offers a practical lever to boost thermal transport in large, suspended graphene, whereas in supported devices mirror-symmetry breaking and substrate scattering reduce ZA lifetimes and can counteract the strain benefit.[98, 103]
Figure 4: (a) Branch-resolved lattice thermal conductivity of suspended monolayer graphene from first-principles BTE. The ZA branch dominates over the full T-range ($\approx 76\%$ at 300 K). Experimental data over 9.7-µm holes has been shown for comparison. Figure taken from Lindsay et al.[13]. (b) Phonon dispersion under uniaxial tensile strain showing linearization of the ZA branch at small $q$ (and corresponding DOS change), which underpins the strain-enhanced $\kappa_{\mathrm{p}h}$ in suspended graphene. Figure taken from Kuang et al.[103].
@@ 5.3 Effect of first–order electron–phonon interactions on the lattice thermal conductivity
At first-order, a single phonon scatters with a single electron, creating a new electron, EP $\rightarrow$ E. This adds an e-ph channel to the phonon collision operator, and in a phonon Boltzmann framework, this shortens the transport lifetime $\tau_{\nu\mathbf{q}}$ that controls $\kappa_{\mathrm{ph}}$.[104] Using the standard decoupled, linearised phonon BTE, the lattice thermal conductivity can be written as
$\kappa_{\mathrm{ph}}=\frac{1}{N_{\mathbf{q}}\,V\,k_{B}T^{2}}\sum_{\nu\mathbf{q}}n^{0}_{\nu\mathbf{q}}\!\left(n^{0}_{\nu\mathbf{q}}{+}1\right)\,(\hbar\omega_{\nu\mathbf{q}})^{2}\,\mathbf{v}_{\nu\mathbf{q}}\!\cdot\!\mathbf{F}_{\nu\mathbf{q}},$ (45)
where $n^{0}_{\nu\mathbf{q}}$ is the Bose factor, $\omega_{\nu\mathbf{q}}$ and $\mathbf{v}_{\nu\mathbf{q}}$ are the phonon frequency and group velocity, and $\mathbf{F}_{\nu\mathbf{q}}$ is the BTE response vector (in the relaxation–time approximation, $\mathbf{F}_{\nu\mathbf{q}}=\tau_{\nu\mathbf{q}}\,\mathbf{v}_{\nu\mathbf{q}}$).[104] According to Matthiessen’s rule, the total mode lifetime is commonly assembled as
@@
where $\tau_{\nu\mathbf{q}}^{\mathrm{iso}}$ is the isotropic scattering lifetime and $\tau_{\nu\mathbf{q}}^{\mathrm{B}}$ is the boundary scattering lifetime. We derive the first–order e–ph rate, Eq. (35), using the DFPT/Wannier e–ph matrix elements $g^{\nu}_{mn}(\mathbf{k},\mathbf{q})$.[48, 99, 105]
@@
In suspended, mirror–symmetric graphene, the one–phonon coupling to flexural ZA modes is symmetry–forbidden at first-order, so the direct e-ph channel acts primarily on the in–plane LA/TA branches.[106] Iterative (decoupled) phonon–BTE calculations nevertheless find a large net reduction of $\kappa_{\mathrm{ph}}$ when the first–order $(\tau_{\nu\mathbf{q}}^{\mathrm{e-ph}})^{-1}$ is included, because frequent LA/TA–electron events feed back—via abundant normal phonon processes—into the transport lifetimes of all branches (including ZA).[104] Quantitatively, turning on the first-order e-ph channel yields a $\sim\!21\%$ drop of $\kappa_{\mathrm{ph}}$ at 300 K and $\sim\!32\%$ at $200$ K, with a non–monotonic carrier–density trend exhibiting a minimum near $n\simeq 4.9\times 10^{14}\,\mathrm{cm^{-2}}$.[104] This can be seen in Fig. 5(a,c). Furthermore, for graphene, the long–wavelength LA/TA e–ph matrix elements admit analytic forms that match first–principles results and govern the temperature/doping dependence of acoustic e–ph scattering across the Bloch–Grüneisen crossover; these forms feed directly into the $\mathbf{q}$– and $\omega$–dependence of $(\tau_{\nu\mathbf{q}}^{\mathrm{e-ph}})^{-1}$ used in Eq. (35) [106] and is shown in Fig. 5(b,d).
@@
First–principles DFPT/EPW workflows applied to other 2D Dirac crystals, such as silicene, explicitly compute $(\tau_{\nu\mathbf{q}}^{\mathrm{e-ph}})^{-1}$ for each branch (ZA/TA/LA) as functions of frequency and carrier density. Adding these mode–resolved e-ph rates to the intrinsic ph–ph terms (Matthiessen assembly) and solving the phonon BTE iteratively demonstrates a sizable suppression of $\kappa_{\mathrm{ph}}$ under doping: the first–order e-ph channel trims the long mean–free–path tail and reduces branch–resolved conductivities in a manner consistent with symmetry and phase–space constraints.[105]
@@
The three takeaways of first-order e-ph interactions in 2D Dirac crystals are: (i) Mechanism: the first-order e–ph channel adds $(\tau_{\nu\mathbf{q}}^{\mathrm{e-ph}})^{-1}$ that shortens transport lifetimes in the BTE integrand; because normal phonon processes are strong in graphene, LA/TA–electron scattering indirectly reduces ZA transport lifetimes as well.[104] (ii) Magnitude: in graphene this yields $\sim$20–30% reductions of $\kappa_{\mathrm{ph}}$ in the 200–300 K range at finite carrier densities.[104] (iii) Generality: comparable, explicitly first–order reductions are found in other 2D crystals (silicene, phosphorene) once $(\tau_{\nu\mathbf{q}}^{\mathrm{e-ph}})^{-1}$ is included and the BTE is solved beyond RTA.[105]
@@
Figure 5: Graphene lattice thermal conductivity with first-order e-ph scattering. Panels (a,c): total lattice thermal conductivity $\kappa_{\mathrm{ph}}$ from an iterative, decoupled phonon-BTE including the first-order e-ph channel. (a) $\kappa_{\mathrm{ph}}(T)$ for several carrier densities $n$ shows a larger e-ph impact at lower $T$; (c) $\kappa_{\mathrm{ph}}(n)$ at fixed $T$ exhibits a non-monotonic trend with a minimum near $n\!\approx\!4.9\times 10^{14}\,\mathrm{cm^{-2}}$. Panels (b,d): mechanism and transport treatment. (b) Branch-resolved $\kappa_{\mathrm{ph}}$ (LA/TA vs. ZA) demonstrates that while the one-phonon ZA coupling is symmetry-forbidden in suspended graphene, frequent LA/TA electron events (together with abundant normal phonon processes) indirectly suppress the ZA transport contribution; (d) Comparison of relaxation-time approximation (RTA) with the fully iterative solution highlights RTA’s underestimate and the importance of normal-process (vertex) corrections. Together, the panels quantify how first-order e-ph scattering shortens transport lifetimes, trims the long-MFP tail, and redistributes heat among branches even when direct ZA–electron coupling vanishes by symmetry. Figure taken from Yang et al.[104].

### v2

@@ 5.1 Electronic screening: Thomas–Fermi vs Lindhard (static/dynamic)
In calculating the e-ph interaction, the Thomas-Fermi approximation—predicated on a large, nearly spherical Fermi surface and ultrashort screening lengths—breaks down in 2D Dirac crystals, where the Fermi “surface” collapses to a point at low carrier density and e-ph coupling is strongly wave–vector selective. In this regime the electronic screening length can be comparable to, or even exceed, the phonon wavelength, and the static Thomas-Fermi response fails to reproduce the sharp enhancement that occurs as the phonon wave vector approaches the $2k_{F}$ nesting condition. Evaluating the full Lindhard polarisation for acoustic phonons yields a dielectric function with a pronounced cusp at $q=2k_{F}$ in the static limit $(\omega=0)$ and a strong dynamic enhancement (divergence in idealised limits) near $q=2k_{F}$ for $\omega\neq 0$. Behaviour entirely missed by Thomas-Fermi[kazemian2023dynamic]. This replaces the oversimplified Thomas-Fermi picture with a $q$- and $\omega$-dependent screening that faithfully captures Dirac-specific e-ph physics.
A convenient form of the Lindhard dielectric function is
$\varepsilon(q,\omega)=1-V(q)\,\Pi(q,\omega),$ (36)
where $V(q)$ is the bare Coulomb interaction (e.g., $V(q)=2\pi e^{2}/\kappa q$ in 2D, $4\pi e^{2}/\kappa q^{2}$ in 3D) and $\Pi(q,\omega)$ is the noninteracting density–density response (Lindhard polarisation), here evaluated on the phonon frequency $\omega=\omega_{\nu\mathbf{q}}$. Two hallmarks follow for acoustic modes that TF cannot capture: (i) a cuspidal feature at $q=2k_{F}$ in the static limit ($\omega=0$), and (ii) a strong dynamic enhancement near $q=2k_{F}$ for $\omega\neq 0$ (regularised in practice by finite temperature, disorder, and intrinsic damping), precisely where phase space is largest (see Fig. 4 for a schematic comparison). These features directly enter the screened e-ph matrix elements,
$g_{mn\nu}^{\mathrm{scr}}(\mathbf{k},\mathbf{q})\;\propto\;\frac{g_{mn\nu}^{(0)}(\mathbf{k},\mathbf{q})}{\varepsilon(q,\omega_{\nu\mathbf{q}})},$ (37)
and thus the phonon self–energy and linewidth[jishi2013feynman, kazemian2023dynamic, allen1972neutron, debernardi2002anharmonic],
[…]
which feed directly into $\kappa_{\mathrm{ph}}$ and coupled e-ph transport.
Practical guidance. For metals or high–density 2D systems with $q\!\ll\!k_{F}$, Thomas-Fermi or static Lindhard screening is often adequate. In low– to moderately doped 2D Dirac crystals, at least the static Lindhard response is required to capture the $2k_{F}$ cusp; whenever the phonon frequency $\omega_{\nu\mathbf{q}}$ is relevant (acoustic branches, finite-$T$ broadening), dynamic Lindhard screening should be used. Finite temperature smooths the static cusp and finite lifetimes regularise the dynamic enhancement, but the strong wave-vector selectivity around $q\!\approx\!2k_{F}$ remains the key qualitative difference from Thomas-Fermi. In the long-wavelength limit $q\!\to\!0$ at fixed (large) $k_{F}$, both Thomas-Fermi and Lindhard recover the familiar Thomas-Fermi-like $1/q$ trend.
Figure 4: (a) The dielectric response function $\chi$ vs $\Psi=q/2\mathrm{k}_{\mathrm{F}}\approx 1$ for different 2D Dirac crystals with different $\mathrm{w_{F}}=c/\mathrm{v}_{\mathrm{F}}$ at T = 0 K. We observe that the dielectric response function diverges near the FSN at the point $\Psi\approx 1$ even for small values of $\mathrm{w_{F}}\neq 0$. The negligible difference in the dielectric response function of various 2D Dirac crystals with different values of $\mathrm{w_{F}}$ shows the generality of our solution for 2D Dirac crystals. (b) Comparing the dielectric response function for the static case $c=0$, designated with $\chi_{0}(\Psi)$, of a 2D Dirac crystal with that of a 1D and 2D Fermion gas at T = 0 K. We observe that $\chi_{0}(\Psi)$ of a 2D Dirac crystal does not follow a constant line like the 2D Fermi gas and it increases reaching a maximum at the FSN point, making the study of 2D Dirac crystals more intricate. Figure taken from Kazemian et al.[kazemian2023dynamic].
@@ 5.2 Branch dependence of e-ph lifetimes and their impact on the thermal conductivity
with bending rigidity $\kappa_{b}$ and areal mass density $\rho$. In two dimensions these dispersions imply that the phonon density of states (DOS) scales as $D_{\mathrm{LA/TA}}(\omega)\!\propto\!\omega$ for linear modes and is approximately constant for the quadratic ZA branch (contrary to a $\omega^{-1/2}$ scaling); this difference is central to the low–frequency phase space in 2D.[Lindsay2014_PRB, Gu2018_RMP, Nika2017_RPP]
@@
$\Gamma_{\nu\mathbf{q}}\equiv\bigl(\tau^{\mathrm{e\text{-}ph}}_{\nu\mathbf{q}}\bigr)^{-1}=\frac{2\pi}{\hbar}\!\sum_{mn\mathbf{k}}\bigl|g^{\nu}_{mn}(\mathbf{k},\mathbf{q})\bigr|^{2}\bigl[f_{n\mathbf{k}}-f_{m,\mathbf{k+q}}\bigr]\,\delta\bigl(\epsilon_{n\mathbf{k}}-\epsilon_{m,\mathbf{k+q}}+\hbar\omega_{\mathbf{q}\nu}\bigr),$ (40)
where the Pauli factor selects states near $E_{\mathrm{F}}$ and $g^{\nu}_{mn}$ are the DFPT/Wannier matrix elements.[Giustino2017_RMP, Sohier2014_PRB] Before specializing the small $\omega$ scalings of the e-ph linewidths, it is useful to recall how the BTE integrand-the product of the phonon density of states $D_{\nu}(\omega)$, the modal heat capacity $C_{\nu}(\omega)$, the squared group velocity $v_{\nu}(\omega)]^{2}$, and the lifetime $\tau_{\nu}(\omega)$—weights each branch in suspended graphene where the first-order ZA e-ph vertex is symmetry-forbidden. Full, iterative solutions of the phonon BTE (including normal processes) predict a ZA-dominated lattice conductivity and a large gap relative to single-mode RTA. This baseline (ph–ph limited, $\sigma_{h}$ intact) is summarized in Fig. 5(a).
First, in-plane LA/TA modes. At long wavelength, graphene’s in–plane coupling is governed by an unscreened gauge–field (shear) interaction, while the scalar deformation potential is strongly screened. First–principles BTE shows that this gauge term controls electronic resistivity and the low–$q$ e–ph matrix elements for LA/TA modes.[Sohier2014_PRB, Park2014_NanoLett] Near charge neutrality the available electronic phase space is small, but for doped samples $N(E_{\mathrm{F}})\!\propto\!|E_{\mathrm{F}}|$ grows linearly with $|E_{\mathrm{F}}|$ (or $k_{F}$) and enhances $\Gamma_{\mathrm{LA/TA}}$ roughly in proportion to $k_{F}$ at fixed $\omega$. Consequently, the LA/TA contribution to the e–ph–limited lattice conductivity decreases with carrier density through the $k_{F}$–dependence of $\tau_{\mathrm{LA/TA}}$, while retaining the weak qq–dependence characteristic of linear modes.[CastroNeto2009_RMP, DasSarma2011_RMP]
At long wavelength the gauge-field (shear) coupling dominates and is only weakly $q$-dependent, whereas the scalar deformation potential is strongly screened.[Sohier2014_PRB, Park2014_NanoLett] In doped graphene this gives the small-$\omega$ scaling
@@
consistent with the linear-$T$ acoustic resistivity of Dirac carriers.[CastroNeto2009_RMP, DasSarma2011_RMP]
Second, flexural ZA modes and mirror symmetry. For freestanding, mirror–symmetric graphene the first–order (one–phonon) e–ph coupling to ZA modes is symmetry–forbidden; ZA scattering of electrons then proceeds via higher–order/two–phonon processes and is comparatively weak.[Mariani2008_PRL] This same mirror selection rule also restricts anharmonic ph–ph scattering of ZA modes, which helps explain why ZA phonons dominate $\kappa_{\mathrm{ph}}$ in pristine suspended graphene.[Lindsay2014_PRB, Lindsay2010_PRB] When mirror symmetry is broken by a substrate, electrostatic gating, static ripples or edges, a linear one–phonon ZA coupling appears with small-$q$ amplitude $\propto q$, giving
@@
This greatly increases the ZA damping rate and reduces the ZA share of $\kappa_{\mathrm{ph}}$. First–principle studies show this mechanism can become mobility/thermal-transport limiting in supported devices and under gating.[Nika2017_RPP, Gunst2017_PRL]
@@
with bending rigidity $\kappa_{b}$ and areal mass density $\rho$. In two dimensions these dispersions imply that the phonon density of states (DOS) scales as $D_{\mathrm{LA/TA}}(\omega)\!\propto\!\omega$ for linear modes and is approximately constant for the quadratic ZA branch (contrary to a $\omega^{-1/2}$ scaling); this difference is central to the low–frequency phase space in 2D.[Gu2018_RMP, Nika2017_RPP, Lindsay2014_PRB]
@@
showing a stronger ultraviolet weighting than LA/TA. In pristine suspended graphene the linear ZA coupling is forbidden, $\Gamma_{\mathrm{ZA}}\!\to\!0$ at one-phonon level, and ZA modes dominate $\kappa_{\mathrm{ph}}$; on substrates or under gating, the activated ZA coupling enhances ZA damping and shifts heat carriage toward in-plane modes.[Nika2017_RPP, Lindsay2014_PRB, Gunst2017_PRL] It should be noted that Eqs. (46)–(47) capture robust trends. Quantitative predictions require the full mode and $\mathbf{q}$ resolved $g^{\nu}_{mn}$, realistic low-$\omega$ cutoffs (finite size/tension for ZA), and the actual Debye (or numerical) upper limit.[Giustino2017_RMP, Lindsay2014_PRB, Sohier2014_PRB]
@@ 5.3 Effect of tensile strain on flexural modes and the thermal conductivity
Applying a uniform uniaxial tensile strain $\epsilon$ couples the bending rigidity $\kappa_{b}$ to an in-plane membrane tension $\sigma=Y\,\epsilon$ (with $Y$ the 2D Young’s modulus). The flexural (ZA) dispersion is renormalized from purely quadratic to[Nika2017_RPP, Lindsay2014_PRB]
@@
Second, higher group velocity and modified scattering. Linearization raises the ZA group velocity from $v_{\mathrm{ZA}}=\partial\omega/\partial q\!\propto\!q$ (bending) to the strain-controlled constant $v_{\epsilon}$ at small $q$. Combined with the DOS change, this alters the three-phonon phase space and lengthens ZA lifetimes in suspended graphene; within the BTE, larger $v^{2}\tau$ for ZA modes increases their contribution to $\kappa_{\mathrm{ph}}$.[Lindsay2014_PRB]
First-principles BTE calculations show that even modest tensile strain $(\approx 1\%)$ linearizes ZA at long wavelengths and increases the intrinsic lattice thermal conductivity of suspended graphene; in the thermodynamic limit, $\kappa_{\mathrm{ph}}$ becomes length-divergent under tension due to the altered ZA scattering landscape.[Lindsay2014_PRB] For finite samples, the net gain depends on size and edge scattering: when the characteristic length exceeds the dominant ZA mean-free paths (tens of microns and beyond), strain yields sizable $\kappa_{\mathrm{ph}}$ enhancements; for smaller membranes or ribbons, boundary scattering can mask the intrinsic increase, leading to weaker strain dependence.[Nika2017_RPP, Kuang2016_IJHMT] Overall, strain-induced flexural hardening offers a practical lever to boost thermal transport in large, suspended graphene, whereas in supported devices mirror-symmetry breaking and substrate scattering reduce ZA lifetimes and can counteract the strain benefit.[Nika2017_RPP, Kuang2016_IJHMT]
Figure 5: (a) Branch-resolved lattice thermal conductivity of suspended monolayer graphene from first-principles BTE. The ZA branch dominates over the full T-range ($\approx 76\%$ at 300 K). Experimental data over 9.7-µm holes has been shown for comparison. Figure taken from Lindsay et al.[Lindsay2014_PRB]. (b) Phonon dispersion under uniaxial tensile strain showing linearization of the ZA branch at small $q$ (and corresponding DOS change), which underpins the strain-enhanced $\kappa_{\mathrm{p}h}$ in suspended graphene. Figure taken from Kuang et al.[Kuang2016_IJHMT].
@@ 5.4 Effect of first–order electron–phonon interactions on the lattice thermal conductivity
At first-order, a single phonon scatters with a single electron, creating a new electron, EP $\rightarrow$ E. This adds an e-ph channel to the phonon collision operator, and in a phonon Boltzmann framework, this shortens the transport lifetime $\tau_{\nu\mathbf{q}}$ that controls $\kappa_{\mathrm{ph}}$.[yang2021indirect] Using the standard decoupled, linearised phonon BTE, the lattice thermal conductivity can be written as
$\kappa_{\mathrm{ph}}=\frac{1}{N_{\mathbf{q}}\,V\,k_{B}T^{2}}\sum_{\nu\mathbf{q}}n^{0}_{\nu\mathbf{q}}\!\left(n^{0}_{\nu\mathbf{q}}{+}1\right)\,(\hbar\omega_{\mathbf{q}\nu})^{2}\,\mathbf{v}_{\nu\mathbf{q}}\!\cdot\!\mathbf{F}_{\nu\mathbf{q}},$ (50)
where $n^{0}_{\nu\mathbf{q}}$ is the Bose factor, $\omega_{\mathbf{q}\nu}$ and $\mathbf{v}_{\nu\mathbf{q}}$ are the phonon frequency and group velocity, and $\mathbf{F}_{\nu\mathbf{q}}$ is the BTE response vector (in the relaxation–time approximation, $\mathbf{F}_{\nu\mathbf{q}}=\tau_{\nu\mathbf{q}}\,\mathbf{v}_{\nu\mathbf{q}}$).[yang2021indirect] According to Matthiessen’s rule, the total mode lifetime is commonly assembled as
@@
where $\tau_{\nu\mathbf{q}}^{\mathrm{iso}}$ is the isotropic scattering lifetime and $\tau_{\nu\mathbf{q}}^{\mathrm{B}}$ is the boundary scattering lifetime. We derive the first–order e–ph rate, Eq. (40), using the DFPT/Wannier e–ph matrix elements $g^{\nu}_{mn}(\mathbf{k},\mathbf{q})$.[Giustino2017_RMP, Sohier2014_PRB, Yue2019_PRB_SiPh]
@@
In suspended, mirror–symmetric graphene, the one–phonon coupling to flexural ZA modes is symmetry–forbidden at first-order, so the direct e-ph channel acts primarily on the in–plane LA/TA branches.[Kaasbjerg2012_PRB_GrapheneAcoustic] Iterative (decoupled) phonon–BTE calculations nevertheless find a large net reduction of $\kappa_{\mathrm{ph}}$ when the first–order $(\tau_{\nu\mathbf{q}}^{\mathrm{e-ph}})^{-1}$ is included, because frequent LA/TA–electron events feed back—via abundant normal phonon processes—into the transport lifetimes of all branches (including ZA).[yang2021indirect] Quantitatively, turning on the first-order e-ph channel yields a $\sim\!21\%$ drop of $\kappa_{\mathrm{ph}}$ at 300 K and $\sim\!32\%$ at $200$ K, with a non–monotonic carrier–density trend exhibiting a minimum near $n\simeq 4.9\times 10^{14}\,\mathrm{cm^{-2}}$.[yang2021indirect] This can be seen in Fig. 6(a,c). Furthermore, for graphene, the long–wavelength LA/TA e–ph matrix elements admit analytic forms that match first–principles results and govern the temperature/doping dependence of acoustic e–ph scattering across the Bloch–Grüneisen crossover; these forms feed directly into the $\mathbf{q}$– and $\omega$–dependence of $(\tau_{\nu\mathbf{q}}^{\mathrm{e-ph}})^{-1}$ used in Eq. (40) [Kaasbjerg2012_PRB_GrapheneAcoustic] and is shown in Fig. 6(b,d).
@@
First–principles DFPT/EPW workflows applied to other 2D Dirac crystals, such as silicene, explicitly compute $(\tau_{\nu\mathbf{q}}^{\mathrm{e-ph}})^{-1}$ for each branch (ZA/TA/LA) as functions of frequency and carrier density. Adding these mode–resolved e-ph rates to the intrinsic ph–ph terms (Matthiessen assembly) and solving the phonon BTE iteratively demonstrates a sizable suppression of $\kappa_{\mathrm{ph}}$ under doping: the first–order e-ph channel trims the long mean–free–path tail and reduces branch–resolved conductivities in a manner consistent with symmetry and phase–space constraints.[Yue2019_PRB_SiPh]
@@
The three takeaways of first-order e-ph interactions in 2D Dirac crystals are: (i) Mechanism: the first-order e–ph channel adds $(\tau_{\nu\mathbf{q}}^{\mathrm{e-ph}})^{-1}$ that shortens transport lifetimes in the BTE integrand; because normal phonon processes are strong in graphene, LA/TA–electron scattering indirectly reduces ZA transport lifetimes as well.[yang2021indirect] (ii) Magnitude: in graphene this yields $\sim$20–30% reductions of $\kappa_{\mathrm{ph}}$ in the 200–300 K range at finite carrier densities.[yang2021indirect] (iii) Generality: comparable, explicitly first–order reductions are found in other 2D crystals (silicene, phosphorene) once $(\tau_{\nu\mathbf{q}}^{\mathrm{e-ph}})^{-1}$ is included and the BTE is solved beyond RTA.[Yue2019_PRB_SiPh]
@@
Figure 6: Graphene lattice thermal conductivity with first-order e-ph scattering. Panels (a,c): total lattice thermal conductivity $\kappa_{\mathrm{ph}}$ from an iterative, decoupled phonon-BTE including the first-order e-ph channel. (a) $\kappa_{\mathrm{ph}}(T)$ for several carrier densities $n$ shows a larger e-ph impact at lower $T$; (c) $\kappa_{\mathrm{ph}}(n)$ at fixed $T$ exhibits a non-monotonic trend with a minimum near $n\!\approx\!4.9\times 10^{14}\,\mathrm{cm^{-2}}$. Panels (b,d): mechanism and transport treatment. (b) Branch-resolved $\kappa_{\mathrm{ph}}$ (LA/TA vs. ZA) demonstrates that while the one-phonon ZA coupling is symmetry-forbidden in suspended graphene, frequent LA/TA electron events (together with abundant normal phonon processes) indirectly suppress the ZA transport contribution; (d) Comparison of relaxation-time approximation (RTA) with the fully iterative solution highlights RTA’s underestimate and the importance of normal-process (vertex) corrections. Together, the panels quantify how first-order e-ph scattering shortens transport lifetimes, trims the long-MFP tail, and redistributes heat among branches even when direct ZA–electron coupling vanishes by symmetry. Figure taken from Yang et al.[yang2021indirect].